16 years from now, Paula’s age will be twice her age 12 years ago. Find her present age.

Answer by mananth(16946)


16 years from now, Paula’s age will be twice her age 12 years ago. Find her present age.
Let Paula's present age be x
16 years hence her age will be x+16
12 years ago her age was x-12
x+16=2(x-12)
x+16 = 2x-24
x=40
Paula's present age 40 years